# SINDI: An Efficient Index for Sparse Vector Approximate Maximum Inner Product Search

### v2

@@ 4.1. Pruning Strategies
As discussed in Section 3, for full-precision Sindi the upper bound for computing the inner product between $\vec{x}_{i}$ and $\vec{q}$ is $\varTheta\!\left(\frac{\|q\|}{s}\right)$. For a given $\vec{q}$, the overall query complexity is $\mathcal{O}\!\left(\frac{\|q\|\,l}{s}\right)$, where $l$ is the average number of inverted lists traversed. Thus, reducing query latency requires decreasing $l$, $\|x\|$, and $\|q\|$. This can be achieved via three approaches: list pruning, document pruning, and query pruning. List and document pruning are performed during index construction, while query pruning is applied at query time. In this work, we focus on the construction stage, since query pruning and document pruning are both forms of vector pruning. We compare three strategies—List Pruning (LP), Vector Number Pruning (VNP), and Mass Ratio Pruning (MRP)—and analyze their respective strengths and weaknesses.
@@
Figure 6. Intuition of Pruning and Reorder
@@
Let $\vec{x}\in\mathbb{R}^{d}$ be a vector. The mass of $\vec{x}$ is defined as the sum of the absolute values of $x$’s non-zero entries:
$mass(\vec{x})\;=\;\sum_{x^{j}\in x}\lvert x^{j}\rvert$.
@@
Let $\vec{x}\in\mathbb{R}^{d}$ and let $\pi$ be a permutation that orders the non-zero entries of $\vec{x}$ by non–increasing absolute value, i.e., $|x^{\pi_{j}}|\geq|x^{\pi_{j+1}}|$. For a constant $\alpha\in(0,1]$, let $1\leq r\leq\|x\|$ be the smallest integer satisfying
$\sum_{j=1}^{r}|x^{\pi_{j}}|\;\geq\;\alpha\times mass(\vec{x}).$
The $\alpha$-mass subvector, denoted $\alpha\text{-}mass(\vec{x})$, is the vector whose non-zero entries are $\{x^{\pi_{j}}\}_{j=1}^{r}$.
Figure 7. An Example of List Pruning, Vector Number Pruning and Mass Ratio Pruning.
@@ 5 $\mathcal{D^{\prime}}\leftarrow\mathcal{D^{\prime}}\,\cup\vec{x}^{\prime}_{i}$
Algorithm 3 outlines the construction of the approximate Sindi index. Given a sparse dataset $\mathcal{D}$ of maximum dimension $d$, window size $\lambda$, and pruning ratio $\alpha$, the algorithm first initializes an empty set $\mathcal{D}^{\prime}$ to store pruned vectors (Line 1). For each vector $\vec{x}_{i}\in\mathcal{D}$ (Line 2), its $\alpha$-mass subvector $\alpha\text{-}mass(\vec{x}_{i})$ is computed and assigned to $\vec{x}^{\prime}_{i}$ (Line 3), which is then added to $\mathcal{D}^{\prime}$. The remaining steps invoke PreciseSindiConstruction on $\mathcal{D}^{\prime}$ (Line 5), followed by returning both the inverted index $I$ and the original dataset $\mathcal{D}$ (Line 6) so that reordering can be performed during retrieval.
@@ 11 $dis^{\prime}\leftarrow 1-\delta(\vec{x}_{i},\vec{q})$
Algorithm 4 shows the search procedure of the approximate Sindi index. Given a query $\vec{q}$, inverted index $I$, pruning ratio $\beta$, reordering size $\gamma$, and target $k$, the algorithm initializes an empty min-heap $H$ for coarse candidates and a max-heap $R$ for the final top-$k$ results (Lines 1–2). It first derives the $\beta$-mass subvector $\vec{q}^{\prime}$ (Line 3) and invokes PreciseSindiSearch to obtain $\gamma$ coarse candidates stored in $H$ (Line 4). While $H$ is not empty (Line 5), the best coarse candidate (i,dis)(i,\,dis) is popped (Line 6), its exact distance to $\vec{q}$ is computed as $dis^{\prime}$ (Line 7), and $(i,\,dis^{\prime})$ is inserted into $R$ if it improves the current set or if $R$ contains fewer than $k$ elements (Line 8). If $R$ exceeds size $k$, its worst element is removed (Lines 9–10). After all candidates are processed, $R$ is returned as the final result (Line 12).
5. Experimental Study
5.1. Experimental Settings
Datasets. Table 3 summarizes the datasets used in our experiments, covering: (i) English datasets from the MsMarco (Nguyen et al., 2016) passage ranking benchmark (including SPLADE-1M and SPLADE-FULL) and the NQ (Kwiatkowski et al., 2019) (Natural Questions) benchmark, all trained with the Splade model; (ii) Chinese dataset AntSparse, real business data from Ant Group trained with the BGE-M3 (Chen et al., 2024) model, which has higher dimensionality due to the larger Chinese vocabulary; (iii) Random datasets with non-zero entry dimensions and values drawn from a uniform distribution. For each dataset, Table 3 reports the average number of non-zero entries per vector ($\mathrm{avg}~\|\vec{x}_{i}\|$), average vectors per inverted list ($\mathrm{avg}~l$), and sparsity. The sparsity of $\mathcal{D}$ is: $\textit{sparsity}=1-\frac{\sum_{\vec{x}\in\mathcal{D}}\|\vec{x}\|}{\|\mathcal{D}\|\cdot d}$.
@@ 5.1. Experimental Settings
  • Seismic (Bruch et al., 2024): A sparse vector index based on inverted lists.
  • SOSIA (Zhao et al., 2024): A sparse vector index using min-hash signatures.
  • BMP (Mallia et al., 2024; Carlson et al., 2025): A dynamic pruning strategy for learning sparse vector retrieval. It divides the original dataset into fine-grained blocks and generates a maximum value vector for each block to evaluate whether the block should be queried.
  • HNSW (Malkov and Yashunin, 2020): A graph-based index originally for dense vectors; we modify the data format and distance computation to support sparse vectors.
  • PyANNS (PyANNS Contributors, 2025): The open-source champion of the BigANN Benchmark 2023 Sparse Track. It is built on HNSW and incorporates quantization, query pruning, and rerank strategies.
Parameter Settings. We use the optimal parameters for each algorithm to ensure a fair comparison. Parameter choices either follow the recommendations from the original authors or are determined via grid search. Details are shown in Table 4.
Table 3. Dataset Statistics and Characteristics
@@
Table 4. Construction Parameter Settings
@@
PyANNS max_degree = 32, ef_construction = 1000
@@
Environment. Experiments are conducted on a server with an Intel Xeon Platinum 8269CY CPU @ 2.50GHz and 512 GB memory. We implement Sindi in C++, compiled with g++ 10.2.1 using -Ofast and AVX-512 instructions.
@@
Figure 8. Overall Performance.
5.2. Overall Performance
5.2.1. Recall and QPS
@@
On the Chinese AntSparse-10M dataset encoded by the BGE-M3 (Chen et al., 2024) model, Sindi also performs best. Fixing Recall@50 at 97%, its QPS is $2.5\times$ that of Seismic, the recall of PyANNS is limited due to the high sparsity of the dataset.
On RANDOM-5M, generated uniformly at random, the uniform distribution yields very few common non-zero dimensions between vectors, causing graph-based methods (PyANNS, HNSW) to suffer severe connectivity loss and low recall. The clustering effectiveness of Seismic is also sensitive to data distributions, resulting in marked degradation. In contrast, Sindi is unaffected by data distribution and achieves the best performance, with QPS exceeding Seismic by an order of magnitude.
@@
5.2.2. Index Size and Construction Time
Figure 9 summarizes the index size and construction time of Sindi, Seismic, and PyANNS on the two largest datasets (SPLADE-FULL and AntSparse-10M), showing that Sindi has the lowest construction cost.
Seismic, which stores summary vectors for each block, yields the largest index size; on AntSparse-10M, its size is $3.8\times$ that of Sindi. By contrast, the graph index construction of PyANNS requires numerous distance computations to find neighbors, resulting in a construction time $71.5\times$ that of Sindi on SPLADE-FULL. In comparison, Sindi mainly sorts each vector’s non-zero entries for pruning, keeping the cost low and enabling rapid index building.
@@ 5.2.2. Index Size and Construction Time
Figure 9. Index Size and Construction Time for Different Datasets and Algorithms.
5.3. Parameters
5.3.1. The Impact of $\alpha$
@@
Figure 10. The Impact of $\alpha$.
This experiment examines how the document pruning parameter $\alpha$, which controls the proportion of high-mass non-zero entries retained, affects Sindi’s performance. A larger $\alpha$ retains more non-zero entries per vector, potentially increasing recall but also raising the search cost. We vary $\alpha$ from 0.4 to 0.8 (step 0.1) on SPLADE-FULL, and from 0.7 to 1.0 (step 0.05) on AntSparse, keeping $\beta$ and $\gamma$ fixed.
@@
The slower recall gain at high $\alpha$ is due to the saturation effect in Section 4.1: once enough non-zero entries are kept, further additions barely reduce inner product error. The sharper QPS drop on AntSparse comes from its lower variance of non-zero values, which leads to more retained entries for the same $\alpha$ increase, and thus more postings to scan.
5.3.2. The Impact of sparsity
@@
Figure 11. QPS of Sindi and Seismic on RANDOM-1M Dataset with Different Sparsity
[…]
5.4. Ablation
5.4.1. The Impact of Pruning Method
@@
Figure 12. Recall@10 vs QPS on MsMarco and AntSparse of Mass Ratio Pruning, List Pruning and Vector Number Pruning.
@@
5.4.2. The Impact of Reorder
@@
Figure 13. Reorder vs. Non-Reorder on SPLADE-FULL and AntSparse-10M Datasets: Time Cost and Recall@50 with Varying $\alpha$.
To investigate the impact of the reordering strategy on the performance of Sindi, we compared the cases with and without reordering on the SPLADE-FULL and AntSparse datasets. The query prune ratio $\beta$ was set to 0.2, and the reordering number $\gamma$ was fixed at 500. For the document prune ratio $\alpha$, we varied it from 0.3 to 0.6 in steps of 0.1 on SPLADE-FULL, and from 0.7 to 1.0 in steps of 0.1 on AntSparse-10M. The non-reordering strategy does not require storing the dataset. We evaluated both strategies using query time and Recall@50 as performance metrics.
5.5. Scalability
() SPLADE-FULL
(a) AntSparse-10M
Figure 14. Multi-threaded Scalability of Sindi (QPS) at Different Recall@50 Targets on SPLADE-FULL and AntSparse-10M Datasets.
@@ 5.5. Scalability
6. Related Work
Existing methods for MIPS on sparse vectors employ various index structures, including inverted index, graph, and hash-based designs. Inverted index-based methods, such as Seismic (Bruch et al., 2024), maintain postings for each dimension and prune low-value entries, often clustering postings into blocks, but each vector may appear in multiple postings, causing random access to scattered memory and reducing locality.
Graph-based methods, such as the HNSW-based PyANNS (Malkov and Yashunin, 2020), organize vectors as nodes in a proximity graph, but vector sparsity weakens connectivity and greedy traversal still incurs frequent random memory accesses.
Hash-based methods, like Sosia (Zhao et al., 2024), transform sparse vectors into sets via the SOS representation and apply min-hash to estimate Jaccard similarity. Multiple hash functions improve accuracy but scatter storage across many buckets, hurting locality and making direct computation inefficient.
BMP (Mallia et al., 2024; Carlson et al., 2025) stores non-zero values in postings and prunes them using block-level maximum value vectors. While this avoids common-dimension lookups, overly fine-grained block partitioning increases the number of candidate blocks to evaluate, lengthening query time. In contrast, our method keeps value-stored postings without incurring such overhead.
7. Conclusion

### v3

@@ IV-A Pruning Strategies
As discussed in Section III, for full-precision Sindi the upper bound for computing the inner product between $\vec{x}_{i}$ and $\vec{q}$ is $\varTheta\!\left(\frac{\|q\|}{s}\right)$. For a given $\vec{q}$, the overall query complexity is $O\!\left(\frac{\|q\|\,l}{s}\right)$, where $l$ is the average number of inverted lists traversed. Thus, reducing query latency requires decreasing $l$, $\|x\|$, and $\|q\|$. This can be achieved via three approaches: list pruning, document pruning, and query pruning. List and document pruning are performed during index construction, while query pruning is applied at query time. In this work, we focus on the construction stage, since query pruning and document pruning are both forms of vector pruning. We compare three strategies—List Pruning (LP), Vector Number Pruning (VNP), and Mass Ratio Pruning (MRP)—and analyze their respective strengths and weaknesses.
As discussed in Section III, for full-precision Sindi the upper bound for computing the inner product between $\vec{x}_{i}$ and $\vec{q}$ is $\varTheta\!\left(\frac{\|q\|}{s}\right)$. For a given $\vec{q}$, the overall query complexity is $O\!\left(\frac{\|q\|\,l}{s}\right)$, where $l$ is the average number of inverted lists traversed. Thus, reducing query latency requires decreasing $l$, $\|x\|$, and $\|q\|$. This can be achieved via retaining entries with the largest magnitude. While magnitude-based pruning is an established concept—exemplified by List Pruning (LP) in Seismic [seismic] and utilized in Bmp [bmp]Sindi’s Mass Ratio Pruning (MRP) introduces a distinct objective. Unlike prior methods that employ pruning primarily to shrink the search space (filtering), MRP is architected to minimize inner product error via an adaptive cumulative mass ratio, ensuring high-quality candidates for reordering. In the following, we systematically compare LP, Vector Number Pruning (VNP), and MRP to analyze their respective trade-offs.
@@
Figure 6: Intuition of Pruning and Reorder
@@
Let $\vec{x}\in\mathbb{R}^{d}$ be a vector. The mass of $\vec{x}$ is defined as the sum of the absolute values of $x$’s non-zero entries: $mass(\vec{x})\;=\;\sum_{x^{j}\in x}\lvert x^{j}\rvert$.
@@
Let $\vec{x}\in\mathbb{R}^{d}$ and let $\pi$ be a permutation that orders the non-zero entries of $\vec{x}$ by non–increasing absolute value, i.e., $|x^{\pi_{j}}|\geq|x^{\pi_{j+1}}|$. For a constant $\alpha\in(0,1]$, let $1\leq r\leq\|x\|$ be the smallest integer satisfying $\sum_{j=1}^{r}|x^{\pi_{j}}|\;\geq\;\alpha\times mass(\vec{x}).$ The $\alpha$-mass subvector [seismic], denoted $\alpha\text{-}mass(\vec{x})$, is the vector whose non-zero entries are $\{x^{\pi_{j}}\}_{j=1}^{r}$.
Figure 7: An Example of List Pruning, Vector Number Pruning and Mass Ratio Pruning.
@@ 5  $\mathcal{D^{\prime}}\leftarrow\mathcal{D^{\prime}}\,\cup\vec{x}^{\prime}_{i}$
Algorithm 3 outlines the construction of the approximate Sindi index. Given a sparse dataset $\mathcal{D}$ of maximum dimension $d$, window size $\lambda$, and pruning ratio $\alpha$, the algorithm first initializes an empty set $\mathcal{D}^{\prime}$ to store pruned vectors (Line 1). For each vector $\vec{x}_{i}\in\mathcal{D}$ (Line 2), its $\alpha$-mass subvector $\alpha\text{-}mass(\vec{x}_{i})$ is computed and assigned to $\vec{x}^{\prime}_{i}$ (Line 3), which is then added to $\mathcal{D}^{\prime}$. The remaining steps invoke PreciseSindiConstruction (Algorithm 1) on $\mathcal{D}^{\prime}$ (Line 5), followed by returning both the inverted index $I$ and the original dataset $\mathcal{D}$ (Line 6) so that reordering can be performed during retrieval.
@@ 11  $dis^{\prime}\leftarrow 1-\delta(\vec{x}_{i},\vec{q})$
Algorithm 4 shows the search procedure of the approximate Sindi index. Given a query $\vec{q}$, inverted index $I$, pruning ratio $\beta$, reordering size $\gamma$, and target $k$, the algorithm initializes an empty min-heap $H$ for coarse candidates and a max-heap $R$ for the final top-$k$ results (Lines 1–2). It first derives the $\beta$-mass subvector $\vec{q}^{\prime}$ (Line 3) and invokes PreciseSindiSearch (Algorithm 2) to obtain $\gamma$ coarse candidates stored in $H$ (Line 4). While $H$ is not empty (Line 5), the best coarse candidate (i,dis)(i,\,dis) is popped (Line 6), its exact distance to $\vec{q}$ is computed as $dis^{\prime}$ (Line 7), and $(i,\,dis^{\prime})$ is inserted into $R$ if it improves the current set or if $R$ contains fewer than $k$ elements (Line 8). If $R$ exceeds size $k$, its worst element is removed (Lines 9–10). After all candidates are processed, $R$ is returned as the final result (Line 12).
V Experimental Study
V-A Experimental Settings
Datasets. Table III summarizes the datasets used in our experiments, covering: (i) English datasets from the MsMarco [msmarco] passage ranking benchmark (including SPLADE-1M and SPLADE-FULL) and the NQ [nq] (Natural Questions) benchmark, all trained with the Splade model; (ii) Chinese dataset AntSparse, real business data from Ant Group trained with the BGE-M3 [bge-m3] model, which has higher dimensionality due to the larger Chinese vocabulary; (iii) Random datasets with non-zero entry dimensions and values drawn from a uniform distribution. For each dataset, Table III reports the average number of non-zero entries per vector ($\mathrm{avg}~\|\vec{x}_{i}\|$), average vectors per inverted list ($\mathrm{avg}~l$), and sparsity. The sparsity of $\mathcal{D}$ is: $\textit{sparsity}=1-\frac{\sum_{\vec{x}\in\mathcal{D}}\|\vec{x}\|}{\|\mathcal{D}\|\cdot d}$.
@@ V-A Experimental Settings
  • Seismic [seismic]: A sparse vector index based on inverted lists.
  • SOSIA [sosia]: A sparse vector index using min-hash signatures.
  • BMP [bmp, sp]: A dynamic pruning strategy for learning sparse vector retrieval. It divides the original dataset into fine-grained blocks and generates a maximum value vector for each block to evaluate whether the block should be queried.
  • HNSW [hnsw]: A graph-based index originally for dense vectors; we modify the data format and distance computation to support sparse vectors.
  • PyANNS [pyanns]: The open-source champion of the BigANN Benchmark 2023 Sparse Track. It is built on HNSW and incorporates quantization, query pruning, and rerank strategies.
  • SHNSW [shnsw]:The open-source runner-up of the BigANN Benchmark 2023 Sparse Track. It is a graph-based index (HNSW variant) that utilizes memory optimization and early termination.
  • SINNAMON [sinnamon]: An inverted index using hashing to compress vectors into dense sketches for SIMD scoring.
Parameter Settings. We use the optimal parameters for each algorithm to ensure a fair comparison, the detailed configuration are in Table IV. Parameter choices either follow the recommendations from the original authors or are determined via grid search, the details are provided in Appendix C.
TABLE III: Dataset Statistics and Characteristics
@@
TABLE IV: Construction Parameter Settings (bold denotes selected optima).
@@
Shnsw SPLADE-1M: $M=16$, efConstruction = 200
SPLADE-FULL: $M=16$, efConstruction = 1000
Sinnamon $m=31$, $h=2$
Hnsw & PyANNS max_degree = 32, ef_construction = 1000
@@
Environment. Experiments are conducted on a server running the CentOS, with an Intel Xeon Platinum 8269CY CPU @ 2.50GHz and 512 GB memory. We implement Sindi in C++, compiled with g++ 10.2.1. For hardware acceleration, Sindi supports both AVX-512 intrinsics and compiler auto-vectorization. To ensure reproducibility and peak performance, the following experiments are conducted using the manual AVX-512 intrinsics implementation. Detailed implementations are provided in Appendix A.
@@
Figure 8: Overall Performance.
V-B Overall Performance
V-B1 Recall and QPS
@@
On the Chinese AntSparse-10M dataset encoded by the BGE-M3 [bge-m3] model, Sindi also performs best. Fixing Recall@50 at 97%, its QPS is $2.5\times$ that of Seismic, the recall of PyANNS is limited due to the high sparsity of the dataset.
On RANDOM-5M, generated uniformly at random, the uniform distribution yields very few common non-zero dimensions between vectors, causing graph-based methods (PyANNS, HNSW, SHNSW) to suffer severe connectivity loss and low recall. The clustering effectiveness of Seismic is also sensitive to data distributions, resulting in marked degradation. In contrast, Sindi is unaffected by data distribution and achieves the best performance, with QPS exceeding Seismic by an order of magnitude.
@@
V-B2 Index Size and Construction Time
Figure 9 summarizes the index size (all including datasets and search index storage) and construction time of Sindi, Seismic, and PyANNS on the two largest datasets (SPLADE-FULL and AntSparse-10M), showing that Sindi has the lowest construction time.
Seismic, which stores summary vectors for each block, yields the largest index size; on AntSparse-10M, its size is $3.8\times$ that of Sindi. By contrast, the graph index construction of PyANNS requires numerous distance computations to find neighbors, resulting in a construction time $71.5\times$ that of Sindi on SPLADE-FULL. In comparison, Sindi mainly sorts each vector’s non-zero entries for pruning, keeping the computational overhead low and enabling rapid index building.
V-C Parameters
V-C1 The Impact of $\alpha$
@@
Figure 9: Index Size and Construction Time for Different Datasets and Algorithms.
@@
Figure 10: The Impact of $\alpha$.
This experiment examines how the document pruning parameter $\alpha$, controlling the retained high-mass non-zero entries, affects Sindi’s performance. A larger $\alpha$ retains more entries per vector, potentially increasing recall but raising search costs. We vary $\alpha$ from 0.4 to 0.8 on SPLADE-FULL and 0.7 to 1.0 on AntSparse, keeping $\beta$ and $\gamma$ fixed.
@@
The slower recall gain at high $\alpha$ is due to the saturation effect in Section IV-A: once enough non-zero entries are kept, further additions barely reduce inner product error. The sharper QPS drop on AntSparse comes from its lower variance of non-zero values, which leads to more retained entries for the same $\alpha$ increase, and thus more postings to scan.
V-C2 The Impact of sparsity
@@
[…]
We also report the sensitivity analysis of reordering depth $\gamma$, which is shown in Appendix E.
V-D Ablation
V-D1 The Impact of Pruning Method
@@
Figure 11: QPS of Sindi and Seismic on RANDOM-1M Dataset with Different Sparsity
@@
Figure 12: Recall@10 vs QPS on MsMarco and AntSparse of Mass Ratio Pruning, List Pruning and Vector Number Pruning.
@@
V-D2 The Impact of Reorder
To investigate the effectiveness of the reordering strategy, we evaluated Sindi’s Recall@50 and query time with and without reordering on the SPLADE-FULL and AntSparse-10M datasets. Regarding parameter selection, $\beta$ and $\gamma$ were fixed; these values serve as representative samples drawn from the optimal intervals identified via grid search. Conversely, to assess robustness across varying index densities, we varied the document prune ratio $\alpha$ from 0.3 to 0.6 on SPLADE-FULL and from 0.7 to 1.0 on AntSparse-10M.
V-E Scalability
@@
Figure 13: Reorder vs. Non-Reorder on SPLADE-FULL and AntSparse-10M Datasets: Time Cost and Recall@50 with Varying $\alpha$.
@@
We also report the performance comparison with Lucene-based systems in Appendix F.
() SPLADE-FULL
(a) AntSparse-10M
Figure 14: Multi-threaded Scalability of Sindi (QPS) at Different Recall@50 Targets on SPLADE-FULL and AntSparse-10M Datasets.
VI Related Work
Inverted Index-based Methods. Algorithms like BMW [bmw], BMP [bmp], and Seismic [seismic] utilize blocking strategies. BMW [bmw]’s dynamic pruning often degenerates to brute force due to smooth weight distributions, causing frequent cache misses. We put the comparisons with BMW in Appendix D. BMP [bmp] improves this by pre-sorting block bounds but incurs prohibitive sorting costs. Seismic [seismic] utilizes geometric blocking for skipping but remains bottlenecked by inefficient exact scoring and random memory access to raw vectors.
Graph-based Methods. PyANNS [pyanns] and SHNSW [shnsw] adapt HNSW [hnsw] for sparse data via quantization and data co-location. However, performance degrades on highly sparse datasets where limited node-query overlap causes greedy routing failures. Furthermore, graph index construction remains significantly more expensive than inverted indices.
Hashing-based Methods. SINNAMON [sinnamon] and SOSIA [sosia] explore hashing pathways. SINNAMON [sinnamon] uses dense sketches for SIMD streaming but lacks static pruning, forcing the evaluation of extensive candidates. SOSIA [sosia] employs MinHash for LSH-based estimation; achieving high accuracy requires numerous hash functions, incurring significant overhead, while its randomized storage forces expensive random memory access.
VII Conclusion